A grass seed company conducts a study to determine the relationship between the density of seeds planted (in pounds per 500 sq ft) and the quality of the resulting lawn. Eight similar plots of land are selected and each is planted with a particular density of seed. One month later the quality of each lawn is rated on a scale of 0 to 100. The regression equation is given below, where x denotes seed density, and y denotes lawn quality. The correlation coefficient r = .600. y = 33.14 + 4.54x If possible, use the information to obtain an estimate of the mean lawn quality for all lawns sown with a seed density of 4.7. If not possible, state why.
Answer #1

We are given the regression equation as

y = 33.14 + 4.54x

y is the lawn quality

x is the seed density

We are provided with seed density = 4.7

Substituting the value of seed density in the equal we predict the value of law quality

y = 33.14 + 4.54*(4.7) = 54.478

an estimate of the mean lawn quality for all lawns sown with a seed density of 4.7 is 54.478